# Accelerated Mirror Descent for Non-Euclidean Star-convex Functions

### v1

@@ 3 Main results
Our method is given in Algorithm 1. It extends the accelerated method of [18], which addresses the case of smoothness ($\kappa=2$) for the Euclidean norm $\left\lVert\cdot\right\rVert_{2}$ known to be $C=1$-regular and strongly convex with respect to itself $(q=2,\psi(\cdot)=\frac{1}{2}\left\lVert\cdot\right\rVert_{2}^{2})$. Similar to their method, ours maintains an auxiliary sequence $(x_{t})_{t\geq 1}$ of iterates, from which it derives a sequence $(x_{t}^{md})_{t\geq 1}$ of gradient query points as well as the output sequence $(x_{t}^{ag})_{t\geq 1}$ of approximate minimizers. In each iteration, the gradient point $x_{t}^{md}$ is found by binary search, which we extend from [18] to handle weak smoothness in Section 4. While this search is not necessary for convex $F$, it appears to be a reasonable trade-off for the relaxation to star-convexity, see also the discussion in Section 5. The iterates $x_{t}$ and approximate minimizers $x_{t}^{ag}$ are each updated in [18] using gradient descent. However, to address a non-Euclidean norm, we instead employ a mirror descent step with respect to $D^{\psi}$ to update $x_{t}$, and a proximal step with respect to $\frac{1}{q}\left\lVert\cdot\right\rVert^{q}$ to update $x_{t}^{ag}$.
@@
Example 3.1 ($p$-norms).
Before presenting our algorithm and its convergence guarantees, we revisit the important example of $p$-norms $\left\lVert\cdot\right\rVert_{p}$, which are regular with $C=1$ by Lemma 2.1. As our distance-generating function, we choose $\psi(x)=\frac{1}{q}\|x\|_{p}^{q}$. The suitable order $q$ depends on $p$. For $p\in(1,2]$, we set $q=2$. This choice ensures that $\psi$ is strongly convex with respect to $\left\lVert\cdot\right\rVert_{p}$ with a rate $\mu=p-1$ (see [4, 21]). For $p>2$, we set $q=p$. This ensures that $\psi$ is $(2^{-\frac{p(p-2)}{p-1}},p)$-uniformly convex with respect to  $\|\cdot\|_{p}$ (see [34, 10]). For both settings, we will show sharp convergence guarantees of our algorithm, as summarized in Table 1.
Example 3.2 (Composite norms).
For composite norms, we can compose distance-generating functions to design new ones. For instance, reconsider the composite norm (2). We know that $\psi_{1}(x_{1:d/2})=\frac{1}{2}\sum_{i=1}^{d/2}x_{i}^{2}$ is $1$-strongly convex with respect to $\|\cdot\|_{2}$ and $\psi_{2}(x_{d/2+1,d})=\frac{1}{1.5}(\sum_{i=d/2+1}^{d}|x_{i}|^{1.5})^{2/1.5}$ is $0.5$-strongly convex with respect to $\|\cdot\|_{1.5}$. Therefore we consider
$\psi(x):=0.5\psi_{1}(x_{1:d/2})+0.5\psi_{2}(x_{d/2+1,d}).$
For all $x,y\in\mathbb{R}^{d}$, it holds
$D_{\psi}(x,y)$ $=0.5\Big{(}\psi_{1}(x_{1:d/2})-\psi_{1}(y_{1:d/2})-\langle\nabla\psi_{1}(y_{1:d/2}),x_{1:d/2}-y_{1:d/2}\rangle\Big{)}$
$+0.5\Big{(}\psi_{2}(x_{(d/2+1):d})-\psi_{2}(y_{(d/2+1):d})-\langle\nabla\psi_{2}(y_{(d/2+1):d}),x_{(d/2+1):d}-y_{(d/2+1):d}\rangle\Big{)}$
$D_{\psi}(x,y)$ $\geq\frac{0.5}{2}\|x_{1:d/2}-y_{1:d/2}\|_{2}^{2}+\frac{0.5\cdot 0.5}{2}\|x_{(d/2+1):d}-y_{(d/2+1):d}\|_{1.5}^{2}\geq 0.5\frac{\|x-y\|_{2\circ 1.5}^{2}}{2},$
therefore $\psi$ is $0.5$-strongly convex with respect to the composite norm $\|\|_{2\circ 1.5}$. In general, the new strongly convex coefficient of the composite norm is the weakest of the two.
Starting point $x_{1}\in\mathbb{R}$, iterations $T\geq 0$, parameters $(C_{t},\epsilon_{t},\alpha_{t},\eta_{t})_{t\geq 1}$
Set $x_{1}^{ag}=x_{1}$
for $1\leq t\leq T$ do
     Binary search (Algorithm 2) with endpoints $(x_{t},x_{t}^{ag})$ and parameters $(C_{t},\epsilon_{t})$ to find $\lambda_{t}\in[0,1]$, $x_{t}^{md}=\lambda_{t}x_{t}^{ag}+(1-\lambda_{t})x_{t}$ such that $\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle+C_{t}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\epsilon_{t}$
     Compute $x_{t+1}=\arg\min_{x\in\mathbb{R}^{d}}\{\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D^{\psi}(x,x_{t})\}$
     Compute $x_{t+1}^{ag}=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\}$
end for
return $x_{t+1}^{ag}$
The algorithm takes as input the parameters $(C_{t},\epsilon_{t},\alpha_{t},\eta_{t})$, which include the step sizes $(\alpha_{t},\eta_{t})$ and the stopping conditions $(C_{t},\epsilon_{t})$ for the binary search. All parameters are polynomial functions of the horizon $t$, and are detailed in the proofs in Appendix C.
As we will see in Theorem 4.1 the total number of values and gradients queried in each binary search step is $\mathcal{O}(\log T)$. In turn, the total number of gradients queried in $T$ rounds is $\mathcal{O}(T\log T)$.
3.1 Convergence result
In this subsection, we provide the convergence result of our algorithm. The proof is in Appendix C with a proof sketch in Section 3.2.
Theorem 3.1.
In the setting of Assumption 1, Algorithm 1 with the tuning below returns $x_{T+1}^{ag}$ in $T$ iterations with precision
$F(x_{T+1}^{ag})-F_{\star}~{}=~{}\mathcal{O}_{q,\kappa}\left(\frac{LC^{q}\tau^{q}}{T^{\frac{\kappa q+\kappa-q}{q}}}\Big{(}\frac{D^{\psi}(x_{\star},x_{1})}{\alpha L}+C^{q_{\ast}}\left(\frac{\alpha L}{\mu}\right)^{\frac{\kappa}{q-\kappa}}\log(T)\Big{)}\right)$
where $\mathcal{O}_{q,\kappa}$ omits absolute constants depending only on $(q,\kappa)$. This is achieved for any $\alpha>0$ by
$\alpha_{t}&:=\frac{\alpha}{t^{\beta}}\\ \eta_{t}&:=\alpha_{t}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t\Big{)}^{q-1}\\ C_{t}&:=C^{q_{\ast}}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\ast}-1}-\frac{1}{\tau}=\frac{1}{\tau}\Big{(}(q-\beta)t-1\Big{)}\\ \epsilon_{t}&:=\frac{\alpha^{\frac{q}{q-\kappa}}C^{q_{\ast}}}{t\eta_{t}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.\end{aligned}\qquad\text{where}\qquad\begin{aligned} \begin{cases}r&:=\frac{q-\kappa}{\kappa}>0\\ M&:=\big{(}\frac{r}{q}\big{)}^{r}\geq 0\\ \beta&=q-\kappa+\frac{q-\kappa}{q}\end{cases}\end{aligned}$
In particular, we obtain the classic Nesterov accelerated $T^{-2}$ rate for $\kappa=q=2$ and $\alpha\leq\mu/L$.
The optimal tuning for $\alpha$ in the preceding result depends on the precise value of $D^{\psi}(x_{\star},x_{1})$, which is unknown a priori. It also depends on $\log T$, while we would rather keep the algorithm any-time. To avoid both issues, we suggest the following practical tuning.
Corollary 3.2.
Suppose a bound $\frac{1}{\mu}D^{\psi}(x_{\star},x_{1})\leq B$ is available. Then the tuning of Theorem 3.1
[…]
Tuning the parameters. The general step-sizes $\alpha_{t}$ and $\eta_{t}$ are polynomial in the iteration number $t$. The main constraint on them is that their exponent needs to be bounded. We choose $\alpha_{t}=\alpha t^{-\beta}$ with $\alpha>0$ a coefficient we set depending on the information of the Bregman divergence $D^{\psi}(x_{\star},x_{1})$. We further have $\eta_{t}\sim t^{1-\frac{3q-\kappa(1+q)}{q}}$ which gives the classic Lipschitz convex setting $\eta_{t}\sim 1/\sqrt{t}$ for $q=2,\kappa=1$ and the Nesterov step size of $\eta_{t}\sim t$ for $q=\kappa=2$. The choice of $C_{t}\sim t$ is dictated by the mirror descent analysis, and $\epsilon_{t}$ is chosen to decrease sufficiently fast such that $\eta_{t}\epsilon_{t}\lesssim 1/t$ sums to $\mathcal{O}(\log T)$. Full details are provided in Appendix C.
Applying Corollary 3.2 to the $p$-norms of Example 3.1 yields the following convergence rates. We consider a problem where both initialisation $x_{1}$ and minimizer $x_{\star}$ are in the $p$-norm ball of radius $R$, i.e.,  $\left\lVert x_{1}\right\rVert_{p},\left\lVert x_{\star}\right\rVert_{p}\leq R$ (so that $\left\lVert x_{1}-x_{\star}\right\rVert_{p}\leq 2R$ by the triangle inequality). Recall that for $p$-norms, we use distance-generating function $\psi(\cdot)=\frac{1}{q}\left\lVert\cdot\right\rVert_{p}^{q}$ with $q=\max\{2,p\}$. We show in Appendix F that also $D^{\psi}(x_{\star},x_{1})\leq 2R^{q}$.
Corollary 3.3.
In the setting of Assumption 1, with the $p$-norm $\left\lVert\cdot\right\rVert_{p}$ for $1<p\leq 2$, Algorithm 1 with $q=2$ returns an $\epsilon$-optimal solution within
$\mathcal{O}_{\kappa}\left(\left(\frac{L\tau^{2}R^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\log\Big{(}\frac{L\tau^{\kappa}R^{\kappa}}{\epsilon}\Big{)}\right)$
evaluations of the function value and gradient.
Corollary 3.4.
In the setting of Assumption 1, with the $p$-norm $\left\lVert\cdot\right\rVert_{p}$ for $p\geq 2$, Algorithm 1 with $q=p$ returns an $\epsilon$-optimal solution within
$\mathcal{O}_{p,\kappa}\left(\Big{(}\frac{L\tau^{p}R^{\kappa}}{\epsilon}\Big{)}^{\frac{p}{\kappa p+\kappa-p}}\log\Big{(}\frac{L\tau^{\kappa}R^{\kappa}}{\epsilon}\Big{)}\right)$
evaluations of the function values and the gradients.
These two corollaries for star-convex functions match the lower bounds proved for convex functions, completing Table 1.
3.2 Proof Sketch
The start of the proof is to use star-convexity to linearize the objective $F$ around the gradient point $x_{t}^{md}$, and to relate its performance to that of the mirror descent (be-the-regularized leader) iterates
$x_{t+1}=\arg\min_{x\in\mathbb{R}^{d}}\{\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D^{\psi}(x,x_{t})\}.$
That is, using the star-convexity of $F$, we obtain an upper bound for $F(x_{t}^{md})-F_{\star}$ (see (11)). Since the gradient is evaluated at $x_{t}^{md}$, it is natural to have first a potential evaluated at $x_{t}^{md}$
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}{}\leq{}$ $D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ (3)
[…]
As a next step, we consider the acceleration step, i.e.,
$x_{t+1}^{ag}=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\}.$
After computing the new accelerated point $x_{t+1}^{ag}$ from the proximal step, we transform the previous objective potential analysis of $F(x_{t}^{md})-F_{\star}$ into an accelerated descent lemma comparing the original potential $F(x_{t}^{md})$ with the two neighbour accelerated potentials $\left(F(x_{t}^{ag}),F(x_{t+1}^{ag})\right)$. The key (see (16)) is to combine weak smoothness of $F$ with the inexact gradient-trick [8, 10] (Lemma B.3)
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}$ $\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ (4)
$+$ $\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+B_{t},$
with $A_{t},B_{t}$ parameters given in (15). The choice of $\lambda_{t}$ found by our binary search Algorithm 2 neutralizes part of that equation, delivering
@@ 3.2 Proof Sketch
In the end, we obtain the result for iteration $t$
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+A_{t-1}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+\eta_{t}\epsilon_{t}+B_{t}.$ (5)
For an arbitrary $\alpha>0$, we show that considering $\alpha_{t}=\frac{\alpha}{t^{q-\kappa+\frac{q-\kappa}{q}}}$ leads to a sharp upper bound. We also verify that the choice of parameters satisfies $A_{t}-A_{t-1}\leq\frac{\eta_{t}}{\tau}$, $\eta_{t}\epsilon_{t}=\frac{\mu}{t}$ and $B_{t}\approx\frac{1}{t}$. To conclude, we apply the telescopic sum.

### v2

@@ 3 Main results
Our algorithm takes as input the parameter schedules $(C_{t},\epsilon_{t},\alpha_{t},\eta_{t})$, with $(\alpha_{t},\eta_{t})$ being the step sizes and $(C_{t},\epsilon_{t})$ being the stopping conditions for the binary search. All tuned parameters are polynomial functions of the horizon $T$, and are given in Theorem 3.1 and Appendix C.
1:Starting point $x_{1}=x_{1}^{ag}\in\mathbb{R}^{d}$, iterations $T\geq 0$, parameters $(C_{t},\epsilon_{t},\alpha_{t},\eta_{t})_{t\geq 1}$
2:For $1\leq t\leq T$ do
3:   Binary search (Alg. 2) with endpoints $(x_{t},x_{t}^{ag})$ and parameters $(C_{t},\epsilon_{t})$ to find $\lambda_{t}\in[0,1]$, $x_{t}^{md}=\lambda_{t}x_{t}^{ag}+(1-\lambda_{t})x_{t}$ such that $\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle+C_{t}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\epsilon_{t}$
4:  Compute $x_{t+1}=\arg\min_{x\in\mathbb{R}^{d}}\{\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D_{\psi}(x,x_{t})\}$
5:   Compute $x_{t+1}^{ag}=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\}$
6:EndFor
7:Return $x_{T+1}^{ag}$
As we will see in Theorem 4.1, the total number of values and gradients queried in each binary search step is $\mathcal{O}(\log T)$. In turn, the total number of gradients queried in $T$ rounds is $\mathcal{O}(T\log T)$. We note that the optimal tuning of step sizes will depend on the number of iterations $T$.
3.1 Convergence result
In this section, we present the convergence result of our algorithm. The full proof is provided in Appendix C, with a proof sketch given in Section 3.2.
Theorem 3.1.
In the setting of Assumption 1 with $\kappa<q$, Algorithm 1 with the tuning below returns $x_{T+1}^{ag}$ after $T$ iterations with precision
$F(x_{T+1}^{ag})-F_{\star}~{}=~{}\mathcal{O}_{q,\kappa}\left(\frac{L\tau^{q}}{T^{\frac{\kappa q+\kappa-q}{q}}}\left(\frac{D_{\psi}(x_{\star},x_{1})}{\alpha L}+\Big{(}\frac{\alpha L}{\mu}\Big{)}^{\frac{\kappa}{q-\kappa}}\log(T)\right)\right)$
where $\mathcal{O}_{q,\kappa}$ omits constants depending only on $(q,\kappa)$. This is achieved for any $\alpha>0$ by
$\alpha_{t}&:=\Big{(}\tau(q-\beta)\Big{)}^{q-\kappa}\frac{\alpha}{t^{\beta}},\quad\quad\eta_{t}:=\alpha_{t}\Big{(}\frac{t}{\tau(q-\beta)}\Big{)}^{q-1}.\\ C_{t}&:=\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\ast}-1}-\frac{1}{\tau}=\frac{1}{\tau}\Big{(}(q-\beta)t-1\Big{)}\\ \epsilon_{t}&:=\frac{\alpha^{\frac{q}{q-\kappa}}}{t\eta_{t}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.\end{aligned}\qquad\text{where}\qquad\begin{aligned} \begin{cases}r&:=\frac{q-\kappa}{\kappa}>0\\ M&:=\big{(}\frac{r}{q}\big{)}^{r}\geq 0\\ \beta&=q-\kappa+\frac{q-\kappa}{q}\end{cases}\end{aligned}$
If $q$ and $\kappa$ are close to $2$, our rate is close to the classic $T^{-2}$ accelerated rate by [NEMIROVSKII198521]. The case $q=\kappa=2$ requires special handling, as the above rate would yield an infinite exponent in the last term. We introduce a refined tuning for this case along with a tighter analysis in Appendix J.
Tuning the parameters. The general step sizes $\alpha_{t}$ and $\eta_{t}$ are polynomial in the iteration number $t$ with the main constraint being that their exponent must remain bounded. Specifically, we set $\alpha_{t}\sim\alpha t^{-\beta}$, where $\alpha>0$ is a coefficient determined based on the Bregman divergence $D_{\psi}(x_{\star},x_{1})$. Additionally, $\eta_{t}$ is chosen as $\eta_{t}\sim t^{1-\frac{3q-\kappa(1+q)}{q}}$, which recovers the classic Lipschitz convex setting $\eta_{t}\sim 1/\sqrt{t}$ for $q=2,\kappa=1$ and the Nesterov step size of $\eta_{t}\sim t$ for $q=\kappa=2$. The choice of $C_{t}\sim t$ follows from the mirror descent analysis, while $\epsilon_{t}$ is chosen to decrease sufficiently fast such that $\eta_{t}\epsilon_{t}\lesssim 1/t$ sums to $\mathcal{O}(\log T)$. Full details are provided in Appendix C.
Tuning the parameter $\alpha$. The optimal tuning for $\alpha$ in the preceding result depends on the radius $D_{\psi}(x_{\star},x_{1})$, which is unknown a priori. It also depends on $\log T$, while we prefer tuning independent of the time horizon $T$. To avoid both issues, we suggest the following practical tuning.
Corollary 3.2.
If a bound $\frac{1}{\mu}D_{\psi}(x_{\star},x_{1})\leq B$ is available, then the tuning of Theorem 3.1 with $\alpha~{}=~{}\frac{\mu}{L}\left(\frac{(q-\kappa)B}{\kappa}\right)^{\frac{q-\kappa}{q}}$ guarantees $F(x_{T+1}^{ag})-F_{\star}~{}=~{}\mathcal{O}_{q,\kappa}\left(L\tau^{q}B^{\frac{\kappa}{q}}\frac{\log(T)}{T^{\frac{\kappa q+\kappa-q}{q}}}\right)$.
In the next sections, we will sketch the proof of Theorem 3.1 and examine the complexity of binary search in Section 4. After that, we will finish the complexity analysis of our algorithm in Section 5.
3.2 Proof Sketch of Theorem 3.1 (Convergence rate of Algorithm 1)
The proof can be decomposed into three steps. As the first step of the proof, we use star-convexity of the objective $F$ to linearize it around the gradient point $x_{t}^{md}$. Together with the mirror descent iterates $x_{t+1}$ from line 4 of Algorithm 1, we obtain the upper bound
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}{}\leq{}$ $D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ (3)
@@ 3.1 Convergence result
As the next step, we consider the unconstrained proximal step in line 5 of Algorithm 1, i.e.,
$x_{t+1}^{ag}=\arg\min_{x\in\mathbb{R}^{d}}~{}\Big{\{}\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\Big{\}}.$
This proximal step leads to the following gap between two neighbouring accelerated potentials $\left(F(x_{t}^{md}),F(x_{t+1}^{ag})\right)$
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}\leq A_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+B_{t}$
with parameters $A_{t},B_{t}$ specified in (15). The term $B_{t}$ is the classical residual after combining the weak smoothness assumption of $F$ with the inexact gradient-trick [CT93, AGJ18, AGL24]. This step was previously known to hold in the convex case [AGJ18] or in the unconstrained Euclidean case [pmlr-v125-hinder20a] for different reasons. For the non-convex non-Euclidean unconstrained case, one needs to examine carefully the regularity of the norm square (see Appendix A).
After computing the new accelerated point $x_{t+1}^{ag}$ from the middle point $x_{t}$, the previous objective bound (3) can be manipulated into an upper bound on the potential gap between two iterations $\left(F(x_{t}^{ag}),F(x_{t+1}^{ag})\right)$.
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}$ $\leq D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ (4)
$+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+B_{t}.$
As the last step, the choice of $\lambda_{t}$ in line 3 of Algorithm 1 ensures that
@@
By showing that our tuning satisfies $A_{t}-A_{t-1}\leq\frac{\eta_{t}}{\tau}$, $\eta_{t}\epsilon_{t}=\frac{\mu}{t}$ and $B_{t}\approx\frac{1}{t}$, we get
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}\leq D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1})+A_{t-1}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+O(1/t),$ (5)
which telescopes to our result. To obtain the desired accelerated convergence rate in Table 1, we still need to account for the complexity of finding $\lambda_{t}$. In Section 4, we show that this additional cost remains only a logarithmic factor, even in the non-Euclidean case.